# Beyond Content Relevance: Evaluating Instruction Following in Retrieval Models

### v1

@@ 3.2 Result Overview
A key observation is that the WISE and SICR scores of the sparse retrieval model BM25 indicates that models relying solely on lexical matching, without any sensitivity to instruction-based retrieval or context-aware instructions, struggle to interpret and act on complex instructions. BM25’s inability to adapt underscores the limitations of traditional sparse retrieval methods for instruction-following tasks.
@@ 4 Analysis
Point-wise reranking vs. List-wise reranking. Point-wise Ranking scores each document independently, evaluating relevance without considering the other documents in the set. Each document is treated in isolation, and the model predicts its relevance to the query. However, list-wise reranking considers the entire list of documents when making ranking decisions. Instead of scoring documents in isolation, the model looks at the relative relationships between documents in the list and ranks them in a way that optimizes the overall ranking order. List-wise ranking (represented by the gray row in Table 3) are generally better suited than point-wise ranking (represented by red row in Table 3) for instruction-following tasks because they take into account the broader context of the query and all relevant documents. When following instructions, particularly complex ones, it is essential to consider the relative importance of different documents to provide a more relevant, cohesive response. A list-wise approach is more likely to properly order documents based on their overall relevance to the instruction. Therefore, listwise rankings have the potential to achieve instruction following across different queries.
@@ 5 Related Work
The development of dense retrieval models has significantly enhanced the semantic understanding and efficiency of retrieval systems. Existing dense models can be categorized into two types based on their architecture: Bidirectional Embedding Models and Decoder-only Embedding Models. Bidirectional Embedding Models are typically base on BERT (Devlin, 2018) or T5 (Raffel et al., 2020) encoders, performing general embedding tasks. Early models that base on BERT or T5 for efficient text embeddings include Sentence BERT (Reimers, 2019), SimCSE (Gao et al., 2021) and Sentence T5 (Ni et al., 2021). To better accommodate the requirements of text embeddings, researchers have pre-trained these encoders using contrastive learning (Izacard et al., 2021; Wang et al., 2022).Furthermore, these models are fine-tuned using various supervised datasets to enhance their performance in retrieval tasks or other downstream applications (Lee et al., 2024b; Li & Li, 2023). Compared to bidirectional embedding models, decoder-only embedding models initially perform relatively poorly in general embedding tasks, primarily due to their limited capacity to comprehensively capture and utilize contextual information (Brown, 2020). However, many researchers have sought to optimize these models’ performance by introducing contrastive learning methods to address their deficiencies in embedding tasks Neelakantan et al. (2022). Currently, researchers have explored not only the use of synthetic data (Wang et al., 2023) but also a hybrid strategy combining real and synthetic data (Meng et al., 2024; BehnamGhader et al., 2024), achieving significant success in text embedding tasks. Collectively, these advancements in contrastive pre-training, model scaling, and leveraging weak supervision and synthetic data significantly propel the field of retrieval.
@@
The notion of relevance often varies among users Mizzaro (1998). Consequently, queries alone may not fully address all users’ information needs Ruthven & Lalmas (2003), while instructions can expand these intentions beyond the scope of the queries. Recent information retrieval research has recognized this and tried to train retrieval models by combining instructions with queries to enhance their instruction-following capabilities. In general, existing instruction-following models can be divided into two categories based on instruction design methods: task-aware and content-aware instruction retrieval models. Asai et al. (2023) first proposed a general retrieval system with task-level instruction, setting specific instructions for different retrieval tasks to query corresponding results. Subsequently, Su et al. (2023) expanded the scope of instructions so that text embeddings can not only retrieve but also classify and diagnose duplicate problems. However, these task-aware instructions are too general and lack the specificity of user instructions in real scenarios. On this basis, other researchers have developed content-level instructions. Oh et al. (2024) set instructions to adapt to query-text pairs based on user background (such as work, hobbies). Zhang et al. (2024) set exclusionary instructions based on the content differences between query results, accounting for users’ exclusionary needs in queries. Similarly, Weller et al. (2024a) set instructions to distinguish query results by combining exclusion and inclusion. However, user intentions in real scenarios involve both the internal attributes of answers (content, audience, language) and the external features of answers (format, length).
@@ Appendix A More Details of InFoSearch
Figure 4 illustrates the distribution of data across six dimensions in the FollowIR, InstructIR, and InfoSearch datasets. The chart highlights the varying proportions of query-document pairs based on dimensions like Audience, Keyword, Language, Length, Source, and Format. Table 4 shows the source of datasets used to collect query-document pairs for each dimension. Additionally, Table 9 to Table 14 provide specific examples from each dimension in the InfoSearch dataset.
@@
Appendix B Prompt templates for constructing InfoSearch dataset
We present the prompt that used to guide ChatGPT for generating and rephrasing task in Table 5 and Table 6.
Table 5: Prompt Templates for Generating
Task Prompt Template
Generate Instruction ### TASK
You are tasked with generating a natural query with an instruction based on the query and the condition provided by the user. You will be provided with a query and a condition and you need to:
1. Rephrase the core query as the first sentence, making it sound like a natural human query without changing its meaning. 2. Create a second sentence that specifies the search restriction. 3. Ensure each sentence is smooth, concise, reasonable, natural, and realistic, mimicking a real human tone.
### INPUT
Core Query: {core query}
Condition: {condition}
### FORMATTING
For the query with an instruction you generated, your output should be in the following format:
“core query”: <str: the core query I give you >
“condition”: <str: the condition I give you >
“query_with_instruction”: <str: the query with instruction you generated >
Generate Core Query ### TASK
You excel at generating queries based on relevant documents. You will receive a document provided by the user, and your task is to create a concise query related to that document. The query must be no longer than 10 words and should summarize the core content of the document.
### INPUT
Document: {document}
### FORMATTING
For the core query you generated, your output should be in the following format:
“core query”: <str: the core query you generated >
Table 6: Prompt Templates for Rephrasing
@@ Appendix B Prompt templates for constructing InfoSearch dataset
Source ### TASK
For a core query, I need documents from a blog, forum post, or news article. I will provide you with a core query, the corresponding document from a news article. Your task is to rewrite the document as blog and forum post content.
@@
### CAUTION 1. For the blog you generated, you cannot use the core query as blog title directly. You need to rephrase it, but do not changed the semantics of this query. Besides, you need to give various information in the line under the title, such as the author, when it was published, the word “Blog”, and the section it belongs to. All the above information must be random. 2. For a forum post, it must be a form of discussion among multiple users. The usernames need to be random rather than use “use1”, “use2” etc.
### INPUT
Core Query: {core query}
Document: {document}
### FORMATTING
Your output should be in the following format:
“blog”: <str: the blog you generated >
“forum”: <str: the forum post you generated >
Audience ### TASK
@@
I will provide you with a core query and its corresponding document. The target audience for this document is experts. Your task is to Rewrite this document to make it easily understandable for laymen.
### CAUTION
1. Keep the semantics of the document intact. 2. Do not use any technical jargon in the rewritten document for layman.
### INPUT
Query: {query}
Document for expert: {expert}
### FORMATTING
Your output should be in the following format:
“query”: <str: the query I give you >
“layman”: <str: the rewritten document for layman you generated >
Table 7: Prompt Template for Reversing
Task Prompt Template
Reverse Instruction ### TASK
@@
Your expertise lies in interpreting and transforming direct instructions into their opposite or negative forms while maintaining clarity and coherence in the transformed instructions. Your task is to reverse the instruction I give you.
### CAUTION
@@
While reversing the instruction, ensure that the new instruction conveys the opposite meaning accurately. Please keep in mind that the transformation should remain clear and easy to understand, avoiding any ambiguity.
### INPUT
Instruction: instruction
### FORMATTING
“reverse instruction”: <str: the instruction your reverse >
Appendix C Prompt for List-wise Reranking Models
TASK Prompt Template
Rank <||||system||||>
<||||user||||>
[1] {passage 1}
[2] {passage 2}
[num] {passage {num}}
Search Query: {query}.
<||||assistant||||>
Model Generation: [9] >[4] >[20] >… >[13]
@@ Appendix C Prompt for List-wise Reranking Models
Appendix D The complete results of evaluating with InFoSearch dataset

### v2

@@ 3.2 Result Overview
The WISE and SICR scores of the sparse retrieval model BM25 indicate that models relying solely on lexical matching, without sensitivity to instruction-based retrieval or context-aware instructions, struggle to interpret and act on complex instructions. BM25’s inability to adapt underscores the limitations of traditional sparse retrieval for instruction-following tasks.
@@ 4 Analysis
Point-wise reranking vs. List-wise reranking. Point-wise reranking evaluates each document independently, predicting relevance without considering other documents. In contrast, list-wise reranking considers the entire set, optimizing the overall ranking order by leveraging relative relationships between documents. As shown in Table 2, list-wise ranking (gray row) generally outperforms point-wise ranking (red row) for instruction-following tasks, as it better captures the broader query context and the relative importance of documents. This makes list-wise reranking more effective in organizing documents to align with complex instructions, improving relevance and coherence across different queries.
@@ 5 Related Work
The development of dense retrieval models has significantly enhanced the semantic understanding and efficiency of retrieval systems. Existing dense models can be categorized into two types based on their architecture: Bidirectional Embedding Models and Decoder-only Embedding Models. Bidirectional Embedding Models are typically base on BERT (Devlin, 2018) or T5 (Raffel et al., 2020) encoders, performing general embedding tasks. Early models that base on BERT or T5 for efficient text embeddings include Sentence BERT (Reimers, 2019), SimCSE (Gao et al., 2021) and Sentence T5 (Ni et al., 2021). To better accommodate the requirements of text embeddings, researchers have pre-trained these encoders using contrastive learning (Izacard et al., 2021; Wang et al., 2022).Furthermore, these models are fine-tuned using various supervised datasets to enhance their performance in retrieval tasks or other downstream applications (Lee et al., 2024b; Li & Li, 2023). Compared to bidirectional embedding models, decoder-only embedding models initially perform relatively poorly in general embedding tasks, primarily due to their limited capacity to comprehensively capture and utilize contextual information (Brown, 2020). However, many researchers have sought to optimize these models’ performance by introducing contrastive learning methods to address their deficiencies in embedding tasks (Neelakantan et al., 2022). Currently, researchers have explored not only the use of synthetic data (Wang et al., 2023) but also a hybrid strategy combining real and synthetic data (Meng et al., 2024; BehnamGhader et al., 2024), achieving significant success in text embedding tasks. Collectively, these advancements in contrastive pre-training, model scaling, and leveraging weak supervision and synthetic data significantly propel the field of retrieval.
@@
The notion of relevance often varies among users (Mizzaro, 1998). Consequently, queries alone may not fully address all users’ information needs (Ruthven & Lalmas, 2003), while instructions can expand these intentions beyond the scope of the queries. Recent information retrieval research has recognized this and tried to train retrieval models by combining instructions with queries to enhance their instruction-following capabilities. In general, existing instruction-following models can be divided into two categories based on instruction design methods: task-aware and content-aware instruction retrieval models. TART first proposed a general retrieval system with task-level instruction, setting specific instructions for different retrieval tasks to query corresponding results (Asai et al., 2023). Subsequently, Instructor expanded the scope of instructions so that text embeddings can not only retrieve but also classify and diagnose duplicate problems (Su et al., 2023). However, these task-aware instructions are too general and lack the specificity of user instructions in real scenarios. On this basis, other researchers have developed content-level instructions. InstructIR set instructions to adapt to query-text pairs based on user background (such as work, hobbies) (Oh et al., 2024). ExcluIR set exclusionary instructions based on the content differences between query results, accounting for users’ exclusionary needs in queries (Zhang et al., 2024). FollowIR set instructions to distinguish query results by combining exclusion and inclusion (Weller et al., 2024a). PIR focuses on the ability of retrievers to recognize and respond to different perspectives in queries (pir). However, real user intentions involve both internal (content, audience, language) and external (format, length) answer attributes. MAIR proposed a large-scale instruction retrieval benchmark covering 126 different information retrieval tasks, but lacks an explicit evaluation of the model’s instruction following ability (mair).
@@ 6 Conclusion
Acknowledgement
We thank EIT and IDT High Performance Computing Center for providing computational resources for this project. This work is supported by 2035 Key Research and Development Program of Ningbo City under Grant No.2024Z127.
@@ Appendix A More Details of InFoSearch
A.1 Condition Determination
The query for each dimension can be diversified and expanded through various conditions, allowing the same query to correspond to different relevant documents under different conditions. Except for the Keyword dimension, the other five dimensions have fixed conditions. The condition of the Keyword dimension is the keyword in the document that is relevant to the query. Therefore, the condition of the Keyword dimension needs to be determined after filtering out the Query-Document (Q-D) pairs. To achieve this, both the query and its corresponding document were input into GPT-4, generating a unique condition for each Q-D pair. However, GPT-4 occasionally selected irrelevant words, such as “and” or “what”, that did not align with the user scenario when generating keyword conditions. To address this, we meticulously crafted prompt templates (Table 3) for condition extraction, ensuring that the conditions were both unique and representative of each document, accurately reflecting the document’s core theme.
Table 3: A template that generates the specific conditions required for the keyword dimension
Prompt for Condition Generation
### TASK ### Your task is to generate a condition that refines a given query in relation to a provided document. The condition should be: 1. Relevant to the document’s core topic – It must align with the central theme or key content of the document. 2. A meaningful constraint on the query – It should introduce a specific aspect, subtopic, or perspective that naturally extends the original query while still being directly supported by the document. 3. Not a generic or arbitrary restriction – The condition must be logically derived from the document’s content and should not be a trivial or overly broad constraint.
### INPUT ### You will receive a query and a document as input: Query: {query} Document: {document}
### FORMATTING ### Condition: <the condition your generated >
A.2 Data Collection
To ensure the queries realistic and reduce the human cost, we consciously integrate conditions when filtering Q-A pairs from existing datasets or web pages. For instance, in the Format dimension, due to the lack of available multi-format Q&A datasets, we selectively extract Q-D pairs from StackOverflow posts. For posts containing code and detailed official documentation responses, we use their titles as queries and the complete responses as documents under the [StackOverflow] condition. The pure code snippets within the answers and references to official documentation are separately extracted and used as documents under the [Code Snippet] and [Manual] conditions, respectively. Table 4 shows the source of datasets used to collect query-document pairs for each dimension.
@@ A.2 Data Collection
A.3 Instruction Generation
The generation of accurate, concise, and natural instructions is crucial. When searching, users tend to express their intentions using simple, naural language, so the generated instructions must remain brief and clear, closely mirroring conversational style. To achieve this, we employed words such as “smooth”, “natural”, and “realistic” in the prompts (see Table 5) to guide GPT-4 in crafting instructions that emphasize not only semantic accuracy but also the emulation of authentic user expressions. Furthermore, a two-sentence structure for the instructions, first rephrasing the query and then appending specific constraints. This structure effectively separates the core query from the conditions, enhancing the diversity of generated instructions. For example, “What are the most effective exercises for losing weight? Please find discussions from forum posts only.” This two-sentence structure ensures logical clarity and semantic coherence.
Table 5: A prompt template for Generating Instruction
Prompt for Instruction Generation
### TASK ### You are tasked with generating a natural query with an instruction based on the query and the condition provided by the user. You will be provided with a query and a condition and you need to: 1. Rephrase the core query as the first sentence, making it sound like a natural human query without changing its meaning. 2. Create a second sentence that specifies the search restriction. 3. Ensure each sentence is smooth, concise, reasonable, natural, and realistic, mimicking a real human tone.
### INPUT ### Core Query: {core query} Condition: {condition}
### FORMATTING ### Core query: <the core query I give you >
Condition: <the condition I give you >
Query_with_Instruction: <the query with instruction you generated >
A.4 Document Rewriting
When the query and relevant documents fail to meet the instruction requirements, we use GPT-4 to rewrite the existing documents to generate relevant documents. The documents that need to be modified are mainly concentrated in the source, length and audience dimensions, so we set specific prompts for these dimensions respectively (see Table 6). In this process, we experimented with directly generating condition-satisfying documents from the query, but these often exhibited redundant expressions and inconsistent formatting. Therefore, we adjusted the existing documents, ensuring they meet the instruction requirements while preserving naturalness and authenticity in the language.
Table 6: Prompt templates for document rewriting
@@ A.4 Document Rewriting
Source ### TASK ### For a core query, I need documents from a blog, forum post, or news article. I will provide you with a core query, the corresponding document from a news article. Your task is to rewrite the document as blog and forum post content. ### CAUTION ### 1. For the blog you generated, you cannot use the core query as blog title directly. You need to rephrase it, but do not change the semantics of this query. Besides, you need to give various information in the line under the title, such as the author, when it was published, the word “Blog”, and the section it belongs to. All the above information must be random. 2. For a forum post, it must be a form of discussion among multiple users. The usernames need to be random rather than use “use1”, “use2” etc. ### INPUT ### Core Query: {core query} Document: {document} ### FORMATTING ### Your output should be in the following format: Blog: <the blog you generated >
Forum: <the forum post you generated >
@@
Audience ### TASK ### I will provide you with a core query and its corresponding document. The target audience for this document is experts. Your task is to Rewrite this document to make it easily understandable for laymen. ### CAUTION ### 1. Keep the semantics of the document intact. 2. Do not use any technical jargon in the rewritten document for layman. ### INPUT ### Query: {query} Document for expert: {expert} ### FORMATTING ### Query: <the query I give you >
Layman: <the rewritten document for layman you generated >
Length ### TASK ### I will provide you with a core query and its corresponding paragraph. Your task is to rewrite this paragraph into a single sentence and an article. ### CAUTION ### 1. Ensure that both the sentence and article retain the original meaning of the paragraph. ### INPUT ### Core Query: {core query} Paragraph: {paragraph} ### FORMATTING ### Your output should be in the following format: Sentence: <the rewritten single sentence that answers the query >
Article: <the multi-paragraph rewritten document that answers the query>
A.5 Instruction Reversal
In real retrieval, we observed that results already ranked highly tend to remain at the top even after instructions are applied. This makes it difficult to determine whether the improvement in ranking is due to the model’s understanding of the instructions or simply a result of detailed keyword and semantic matching. To address this, we validate the model’s comprehension of the instructions by reversing the semantic meaning of the instructions. For example, “Please answer in Chinese” is reversed to “Please do not answer in Chinese.”
Table 7: A prompt Template for Instruction Reversing
Prompt for Instruction Reversing
@@ A.5 Instruction Reversal
### TASK ### Your expertise lies in interpreting and transforming direct instructions into their opposite or negative forms while maintaining clarity and coherence in the transformed instructions. Your task is to reverse the instruction I give you.
@@
### CAUTION ### While reversing the instruction, ensure that the new instruction conveys the opposite meaning accurately. Please keep in mind that the transformation should remain clear and easy to understand, avoiding any ambiguity.
### INPUT ### Instruction: {instruction}
### FORMATTING ### Reverse Instruction: <the instruction your reverse >
A.6 Hard Negative Generations
While positive documents for the same query under varying conditions may act as negative examples for one another (instruction negatives), we still need to prevent the model from relying solely on prominent features for simple retrieval, thereby neglecting the subtle relationships between the query and the documents. To address this, we use GPT-4 to generate documents that appear to be related to the query topic on the surface but cannot actually answer the query, serving as hard negative documents (query negatives).
Table 8: A prompt templates for generating hard negative
Prompt for Hard Negative Generation
### TASK ### You are tasked with generating a hard negative document based on a given query. A hard negative document should appear superficially relevant to the query but contain critical inaccuracies, misleading details, or subtle contradictions. Follow these steps: 1. Understand the core intent of the query and identify key entities, relationships, or requirements. 2. Generate a document that incorporates some keywords from the query but does not provide a direct or indirect answer to the query. The document should maintain a plausible structure and stay on a related topic while ensuring that no information within it can be used to infer or construct a correct response to the query. 3. Ensure the document is coherent, natural, and realistic, mimicking a genuine but incorrect response.
### INPUT ### Core Query: {core query}
### FORMATTING ### Core query: <the core query I give you >
Hard negative document: <generated document >
A.7 Manual Review
We filtered out anomalous documents from the outputs of 12 retrieval models, selecting those that failed to rank within the top 50 in six or more models or had a relevance score below 0.5 for the query, followed by manual screening. This process aimed to eliminate mislabeled Q-D pairs selected from other datasets or documents inaccurately retrieved through manual search. For these mismatched documents, we proceed with manual replacement. After multiple rounds of screening to ensure the quality of the InfoSearch dataset, the statistical results are summarized in Table 9.
Dimension # $Q$ Avg $|Q|$ $I$ Avg $|I|$ $R$ Avg $|R|$ $D$
Audience 100 9.02 210 20.46 210 15.91 840
Keyword 100 6.30 288 17.90 288 18.92 1152
Format 100 9.16 300 16.65 300 19.31 1200
Language 100 8.75 200 14.09 200 15.74 800
Length 100 8.52 300 15.94 300 16.26 1200
Source 100 7.38 300 18.19 300 15.58 1200
Total 600 1598 1598 6392
To make the data more intuitive, Table 10 to Table 15 provide specific examples from each dimension in the InfoSearch dataset.
@@ A.7 Manual Review
Moreover, Figure 4 illustrates the distribution of data across six dimensions in the FollowIR, InstructIR, and InfoSearch datasets. The chart highlights the varying proportions of query-document pairs based on dimensions like Audience, Keyword, Language, Length, Source, and Format.
Appendix B Evaluation Metrics Analysis
[…]
$min\mbox{-nDCG}(q_{i})=\min_{j\in(1,2,...,m_{i})}s_{ij}$ (7)
$Robustness@k=\frac{1}{n}\sum_{i=1}^{n}min\mbox{-nDCG}(q_{i})$ (8)
[…]
Appendix C Prompt for List-wise Reranking Models
TASK Prompt Template
Rank <||||system||||>
<||||user||||>
[1] {passage 1}
[2] {passage 2}
[num] {passage {num}}
Search Query: {query}.
<||||assistant||||>
Model Generation: [9] >[4] >[20] >… >[13]
Appendix D Sampling of WISE Score Reward Component
Figure 5: Heatmap of the rewards component
Appendix E The complete results of evaluating with InFoSearch dataset